If you roll this dice 12 times what is the probability that you will roll three 6s? Work out both by hand for the Binomial distribution.
1
Expert's answer
2019-10-07T09:32:05-0400

Take three different numbers "a,b,c" between "1" and "12". The probability of obtaining 6 on dice is "\frac{1}{6}", probability of obtaining other number is "1-\frac{1}{6}=\frac{5}{6}" . So probability of obtaining "6" for "a"-th, "b"-th and "c"-th times and obtaining other numbers for other times is "\left(\frac{1}{6}\right)^3\left(\frac{5}{6}\right)^{12-3}=\frac{5^9}{6^{12}}".

The number of choosing "a, b" and "c" is "\binom{12}{3}", so the probability of obtaining three 6s is "\binom{12}{3}\frac{5^9}{6^{12}}"

Answer: "\binom{12}{3}\frac{5^9}{6^{12}}"